OpenStudy (anonymous):

will fan and medal!!!!!!!!!!!!!! The depth of the water at the end of a pier changes periodically along with the movement of tides. On a particular day, low tides occur at 12:00 a.m. and 3:30 p.m., with a depth of 3.25 meters, while high tides occur at 7:45 a.m. and 11:15 p.m., with a depth of 8.75 meters. Which of the following equations models d, the depth of the water in meters, as a function of time, t, in hours? Let t = 0 be 12:00 a.m.

OpenStudy (anonymous):

answer choices- A. d=-3.75cos(4pi/31 t)+5 B. d=-3.75cos(4pi/29 t)+5 C. d=-2.75cos(4pi/31 t)+6 D. d=-2.75cos(4pi/31 t)+6

OpenStudy (anonymous):

@mathmale

OpenStudy (anonymous):

@XhawaiiangirlX

OpenStudy (anaise):

Still there?

OpenStudy (anonymous):

@Anaise do u know how to solve it?